OpenStudy (anonymous):

Find the vertex, focus, and directrix of the parabola. (y+2)^2 = -2(x-6) I found the vertex which is (6,-2) AND 4p=-2 p=-2/4 p=-1/2

OpenStudy (amistre64):

v = (6,-2) is right 4a = -2 a = |-1/2| = 1/2

OpenStudy (amistre64):

it opens to the right so the focus is (6,-1.5)

OpenStudy (amistre64):

and the D: x = -2.5

OpenStudy (amistre64):

thats backwards lol

OpenStudy (amistre64):

it opens to the left; becasue of the -2.... switch the focus and directix i gave yuo

OpenStudy (anonymous):

oh.. ok

OpenStudy (anonymous):

the focus is wrong. says the program

OpenStudy (amistre64):

focus should be...doh!!, modify the x not the y grrr!!

OpenStudy (amistre64):

V = (6,-2) F=(5.5,-2) D= x=6.5